A person can hear sound waves in the frequency range 20 Hz to 20 kHz. Find the minimum and the
maximum wavelengths of sound that is audible to the person. The speed of sound is 360 m/s.
 

Grade:12

1 Answers

Arun
6 years ago
for maximum wavelength n = 20 Hz. As (η∝1/λ) b) for minimum wavelength, n = 20 kHz ∴ λ = 360/ (20 × 10^3) = 18 × 10^3 m = 18 mm ⇒ * = (v/n) = 360/20 = 18 m